### v1

# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to source material remains a significant challenge due to the generation of hallucinations. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucination within generated text, particularly in long outputs. In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization
[…]
1 Introduction
Recent advancements in Large Language Models (LLMs) have pushed the boundaries of human-like language generation, particularly in tasks like text summarization Chang et al. (2024b). With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Roit et al. (2023); Song et al. (2025). This capability is transforming how we interact with large volumes of text, making information more accessible Achiam et al. (2023); Grattafiori et al. (2024); Team et al. (2024); GLM et al. (2024).
Figure 1: Comparison of the faithfulness scores of summaries generated around 50 words, 200 words, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
As LLMs continue to evolve, a paradigm shift is emerging from short-form to long-form generation, enabled by the ability to process extended contexts Wu et al. (2025a). Long-output generation is critical for complex reasoning tasks like long Chain-of-Thought prompting Jaech et al. (2024) and language agents Sumers et al. (2024); Wang et al. (2024), inherently requiring coherent, contextually grounded long responses. While recent benchmarks evaluate long-generation capabilities Wu et al. (2025b); Ye et al. (2025), they often lack contextual grounding, overlooking a crucial aspect of long-form generation: context-aware faithfulness.
A persistent and critical challenge faced by LLMs is the phenomenon of hallucination, wherein the generated content is unfaithful to or unsupported by the input context Huang et al. (2025). While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the spatial distribution of factual errors within the generated sequence. Related research has explored positional biases in how models process input contexts, notably the Lost in the Middle phenomenon Liu et al. (2024b); Wan et al. (2025). However, understanding the distribution of errors within the generated output itself, particularly for long responses from long contexts, is equally crucial for effective diagnosis and mitigation.
Motivated by this critical research gap, our work provides the first dedicated investigation into the spatial distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization. Unlike prior work primarily addressing long input contexts or short outputs Tang et al. (2023); Zhang et al. (2024a); Bishop et al. (2024), we tackle long context-aware long response generation, which requires processing lengthy inputs and maintaining faithfulness over a significantly long output sequence. Specifically, we pose a core research question: How frequently do hallucinations occur in long-form generation tasks such as document summarization, and when they do, where in the output are they most likely to appear?
Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness (See Section 3.1 for evaluation metric.) significantly decreases towards the end for long summaries (e.g., 800 words). Contrary to expectations of uniform distribution, this distinct positional bias highlights a critical vulnerability in LLMs when generating extended text. To systematically analyze this phenomenon and its implications, our research addresses three core questions:
  1. RQ1.
    Where Do Hallucinations Most Frequently Occur? (§3)
  2. RQ2.
    What Factors Contribute to Hallucination Concentration at the Last Part? (§4)
  3. RQ3.
    How Can We Mitigate the Hallucination in the Last Part? (§5)
To answer these questions, we first empirically characterize the positional distribution of hallucinations (RQ1). Building on this, we explore potential causes of this phenomenon related to the generative process (RQ2). Finally, we propose and evaluate mitigation strategies targeting late-stage hallucinations (RQ3). Our findings highlight the necessity of considering the spatial dimension of errors and provide insights for developing more robust long-form generation systems.
Our contributions can be summarized as follows:
  • We provide the first empirical characterization of the Hallucinate at the Last phenomenon in LLM-based long response generation, particularly in long document summarization.
  • We offer initial insights and analysis into the potential factors contributing to this late-stage hallucination.
  • We investigate mitigation strategies specifically tailored to address positional hallucination.
2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents, varying both the context length and the output length.
2.1 Input Context Lengths
To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 16K tokens. The specific context lengths used in our experiments range from 1K to 8K tokens, with increments of 1K.
2.2 Output Length Categories
We define two categories for output length to study the Hallucinate at the Last phenomenon:
  • Short Output: Summaries with a length between 100 and 200 words.
  • Long Output: Summaries with a length up to 30% of the input context length.
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both short and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix A for more details.
2.3 Overall Faithfulness
To provide a general understanding of the faithfulness levels observed in the generated long summaries before conducting a detailed positional analysis, we first report the overall faithfulness scores.
Context Length Faithfulness(%)
1K 83.3
2K 86.7
3K 78.9
4K 82.9
5K 86.6
6K 85.9
7K 83.9
8K 86.8
Table 1: Faithfulness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
As shown in Table 1, we report the overall faithfulness of long summaries generated by Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset, with context lengths ranging from 1K to 8K. We employ the FineSurE Song et al. (2024) framework with GPT-4 Achiam et al. (2023) to compute the overall faithfulness scores. FineSurE evaluates summaries by comparing each sentence with the source context and identifying specific error types when inconsistencies are detected.
As shown in the results, the overall performance is consistent with the results on the CNNDM Hermann et al. (2015) dataset reported in the original FineSurE paper, where the faithfulness of summaries generated by the Llama3-70B-Instruct model is reported to be 85.5%. Building on this, We perform a more fine-grained evaluation of summary faithfulness in Section 3, which includes a positional analysis to examine how factual consistency varies across different parts of the output.
3 Where Do Hallucinations Most Frequently Occur?
In this section, we empirically investigate where hallucinations most frequently occur in long document summarization. We analyze the positional distribution of factual errors by examining faithfulness across varying models, datasets, and output positions.
Figure 2: Comparison of faithfulness scores for short summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
3.1 Experimental Setup
Models
We conduct experiments using several state-of-the-art Large Language Models to ensure our findings are not model-specific. The models evaluated include Llama3.1-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 Albert Q. Jiang et al. (2023), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini Achiam et al. (2023).
Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
Datasets
To examine whether the phenomenon is domain-specific, we analyze the spatial distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation.
3.2 Evaluation Metric
Faithfulness
We adopt an evaluation metric based on atomic facts, which has shown a significant effectiveness in evaluating summary factuality Min et al. (2023); Tang et al. (2024); Scirè et al. (2024); Jing et al. (2024); Wei et al. (2024); Yu et al. (2024).
Let $s_{i}$ be the $i$-th sentence in the generated summary $S$. Each sentence in the generated summary is decomposed into a set of atomic facts $A_{i}$ by utilizing LLM, followed by a filtering process to remove unnecessary atomic facts Yang et al. (2024b), which makes:
$A_{i}=\{a_{i1},a_{i2},...,a_{iNi}\}$ (1)
Using an NLI model, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner. We measure only the entailment score, and each atomic fact is assigned the highest entailment score among its comparisons:
$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\textrm{{Entail}}(a_{ij},d_{m})$ (2)
Finally, the average of the entailment scores for a set of filtered atomic facts in a sentence is used as the sentence’s faithfulness score:
$\textrm{Faithfulness}(s_{i})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\textrm{score}(a_{ij})$ (3)
See Appendix B for more experimental details.
Sensitivity
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Specifically, each summary is divided into five equal length bins, and the average faithfulness score of the atomic facts within each bin is computed. We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins. A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise. The larger the sensitivity, the more pronounced the hallucination at the end of the output.
3.3 Hallucinate at the Last
The latter part of the summary exhibits the lowest level of faithfulness across nearly all models, and across all contexts and summary lengths.
We present the positional faithfulness discrepancies of both short and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively. As shown in Figure 2, short summaries generated by the three models—excluding Mistral—consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end. Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length. Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries, as shown in Figure 3. As the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below 0.6 when 45 sentences are generated from an 8K context.
We report the sensitivity results for long summaries generated by the Llama3.1 model in Table 2, and those for the Mistral model in Table 3. As shown in the results, summaries generated by Llama3.1 tend to exhibit the lowest faithfulness scores in the final bin, resulting in high sensitivity. In contrast, the Mistral model often shows the lowest faithfulness scores in the initial bins, with sensitivity frequently falling below zero. We analyze the underlying causes of these contrasting patterns in Section 4. See Appendix C for more results.
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.77 0.74 0.69 0.69 0.69 3.3
2K 0.74 0.72 0.70 0.66 0.63 7.5
3K 0.75 0.70 0.69 0.65 0.58 11.8
4K 0.74 0.71 0.68 0.69 0.64 6.5
5K 0.73 0.72 0.72 0.68 0.60 11.3
6K 0.74 0.72 0.66 0.65 0.61 8.3
7K 0.73 0.70 0.69 0.66 0.62 7.5
8K 0.76 0.70 0.72 0.69 0.61 10.8
Table 2: Comparison of faithfulness scores and Sensitivity for long summaries generated by Llama3.1-8B-Instruct across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
2K 0.75 0.80 0.84 0.82 0.81 -0.8
3K 0.81 0.80 0.82 0.84 0.83 -1.3
4K 0.79 0.81 0.83 0.81 0.81 0.0
5K 0.80 0.84 0.86 0.85 0.79 4.8
6K 0.82 0.87 0.87 0.87 0.86 -0.3
7K 0.82 0.84 0.82 0.83 0.85 -2.3
8K 0.87 0.83 0.87 0.89 0.81 5.5
Table 3: Comparison of faithfulness scores and Sensitivity for long summaries generated by Mistral-7B-Instruct-v0.3 across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, contexts, and summary lengths.
We present the positional faithfulness discrepancy of long summaries generated by the Llama3.1-8B-Instruct model across different datasets in Figure 4. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores tend to decrease toward the final segments as the output length increases. In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline. Notably, for the Pubmed dataset, the faithfulness score falls below 0.5 when the model generates 45 sentences at a context length of 8K. Additional sensitivity results across diverse datasets are reported in Appendix D. Notably, across all datasets, none of the computed sensitivity scores fall below zero.
Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K. We report results in short summaries in Appendix D.
4 Why Do Hallucinations Frequently Occur at the Last?
This section investigates the underlying factors that contribute to the Hallucinate at the Last phenomenon, exploring two main hypotheses. The first attributes the phenomenon to the inherent nature of summarization: key information is typically concentrated at the beginning of the summary, with less important content appearing toward the end. The second posits that as LLMs generate longer outputs, they increasingly attend to previously generated tokens rather than the original input context, resulting in a shift in attention distribution.
4.1 Is it Intrinsic to Summarization?
This hypothesis is based on the intrinsic structure of human-written and model-generated summaries. In most summarization tasks—especially in news, scientific, and Wikipedia-style documents—the main ideas and most salient information are typically presented early in the summary Hermann et al. (2015); Narayan et al. (2018); Cohan et al. (2018). This structure mirrors the lead bias often observed in source documents, where the beginning contains the most crucial content Kim et al. (2019); Zhao et al. (2022); Ravaut et al. (2024). Motivated by this, our study investigates whether such structural bias also manifests in the form of decreasing factual consistency in generated outputs. Specifically, we compare the positional faithfulness scores of human-written reference summaries and model-generated summaries to assess whether the decline in faithfulness is an artifact of summarization’s inherent structure or a model-specific behavior.
Figure 5: Comparison of faithfulness scores between the reference summary and summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths on the CNNDM dataset.
Experimental Setup
We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens. We compare these references to summaries generated by the Llama3.1-8B-Instruct model, evaluated using the metric described in Section 3.1.
Results & Analysis
Figure 5 presents a positional faithfulness comparison between human-written reference summaries and model-generated summaries from the CNNDM dataset, analyzed across increasing output lengths. As shown in the results, the reference summaries exhibit a different pattern compared to the generated summaries. While the faithfulness scores of the generated summaries drop below 0.6 in the final segment, the reference summaries show a temporary dip around the middle but recover toward the end. This observation suggests that the Hallucination at the Last phenomenon cannot be solely attributed to the inherent characteristics of summarization.
(a) Average attention weights on the first, middle, and last sentences of summaries generated by the Llama3.1-8B-Instruct model.
(b) Average attention weights on the first, middle, and last sentences of summaries generated by the Mistral-7B-Instruct-v0.3 model.
Figure 6: Average attention weights of two models exhibiting contrasting trends in Figure 3. The dashed line separates the input context from the generated output.
4.2 Is it Intrinsic to Attention?
Prior research on LLMs has demonstrated that attention weight distributions are closely correlated with the generation process, influencing output coherence Dong et al. (2021); Zhang et al. (2024b). Recent studies on Large Vision-Language Models (LVLMs) further suggest that hallucinations frequently occur in the later parts of generated text Liu et al. (2024a); Lee et al. (2024); Min et al. (2025). This phenomenon has been attributed to a shift in attention: as text generation progresses, attention weights increasingly favor previously generated text tokens over image tokens. Inspired by this observation, we investigate whether a similar trend exists in LLMs by analyzing how attention weights on generated tokens evolve as output length increases.
Experimental Setup
Unlike previous studies that focus on attention to the first generated token Hsieh et al. (2024), we compute attention weights at the sentence level. Moreover, we segment the full sequence into chunks of 100 tokens. Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence. See Appendix E for more details.
Results & Analysis
Figure 6 presents a visualization of average attention weights across sentences. As shown in the results, the Llama3.1-8B-Instruct model—which exhibited the Hallucinate at the Last pattern in Figure 3—assigns nearly four times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Mistral-7B-Instruct-v0.3 model assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination. Moreover, we observe that for most models—excluding Mistral—as output length increases, attention becomes increasingly concentrated on generated tokens, which in turn amplifies hallucination.
5 How to Mitigate the Hallucination at the Last?
We investigate how we can resolve the Hallucinate at the Last phenomenon. To this end, we apply four methods.
Experimental Setup
In this experiment, we generate summaries for the Wikipedia dataset using four different methods, with a context length of 7K tokens. As a baseline, we use summaries generated by the Llama3.1-8B-Instruct model. For comparison, we evaluate the following four methods:
  • BooookScore Chang et al. (2024a) segments the input context into chunks, generates summaries for each chunk individually, and then merges the partial summaries.
  • MInference Jiang et al. (2024) employs sparse attention mechanisms to efficiently process long input sequences.
  • LongWriter-Llama3.1-8B Bai et al. (2025) is a model fine-tuned on a long-output dataset and further enhanced using DPO Rafailov et al. (2023).
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding.
See Appendix F for more experimental details.
Methods Generated Summary Bins Sensitivity
1 2 3 4 5
Llama3.1-8B 0.74 0.75 0.75 0.73 0.64 10.3
+ BooookScore 0.73 0.77 0.78 0.80 0.75 2.0
+ MInference 0.77 0.78 0.78 0.71 0.69 7.0
+ LongWriter 0.85 0.77 0.81 0.63 0.70 6.5
+ AdaCAD 0.79 0.82 0.77 0.81 0.64 15.8
Table 4: Comparison of faithfulness scores and Sensitivity for long summaries generated by different mitigation methods across output bins on the Wikipedia dataset with a 7K context length.
Results & Analysis
We report the faithfulness scores across output bins and the corresponding Sensitivity results in Table 4, and the faithfulness scores across increasing output lengths for each method in Figure 7. As shown in Table 4, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. However, while BooookScore mitigates the phenomenon, it doesn’t fundamentally solve the problem of generating long responses directly from raw long inputs, relying instead on decomposing the problem via chunking and merging. This highlights the necessity of future work to develop methods that enable models to maintain faithfulness when generating long text directly from long contexts.
6 Related Work
The problem of hallucination in LLMs has been a significant area of research. Prior work has largely focused on developing methods for detecting factual inconsistencies in generated text Chuang et al. (2024); Hu et al. (2024); Kim et al. (2024); Zhong and Litman (2025) and proposing strategies to mitigate their occurrence, often through improved training Zhang et al. (2023); Wan et al. (2023), decoding Shi et al. (2024); Wang et al. (2025), or prompting Zhou et al. (2023). While effective in reducing overall hallucination rates, these studies typically do not analyze the within-sequence distribution of errors.
The concept of positional bias in generation errors has been observed in related domains. Notably, studies on Large Vision-Language Models (LVLMs) have indicated that hallucinations tend to increase towards the end of generated textual descriptions of images Liu et al. (2024a); Lee et al. (2024); Min et al. (2025).
Recently, several studies have begun to explore long-output generation in LLMs Wu et al. (2025b); Ye et al. (2025). However, these works focus on procedural generation tasks that are not grounded in contextual input, limiting their applicability to real-world summarization scenarios. A notable study on multi-document summarization aligns with our findings Belém et al. (2025). Nevertheless, it does not address positional context-aware faithfulness over extended summary lengths.
Our work presents the first dedicated study of where hallucinations occur in LLM-based long document summarization, moving beyond detection to uncover positional patterns, underlying causes, and mitigation strategies.
Figure 7: Comparison of faithfulness scores for long summaries generated by different mitigation methods across increasing output lengths, using the Wikipedia dataset with a 7K context length.
7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into spatially-aware generation techniques.
Limitations
In this study, we explored four domains, primarily because the corresponding datasets provided input contexts of the desired length. However, we were not able to investigate domains such as books, dialogues, movie scripts, or meeting transcripts. Books were excluded due to their excessive length, while the other domains lacked datasets with suitable input lengths or sufficient sample sizes. Additionally, we also did not explore models of varying sizes.
Despite these limitations, we believe this work offers an important foundation for future research in the relatively underexplored domain of long-form output generation, particularly in summarization.
Ethics Statement
This study leverages publicly available datasets, including Wikipedia, Arxiv, Pubmed, Govreport, and CNNDM, to analyze long-form text generation in LLMs. All experiments were conducted in a consistent and reproducible manner across models and datasets, without any manipulation or omission of data or results. We investigate the phenomenon of Hallucinate at the Last, a tendency observed in long context-aware summarization models to generate hallucinated content toward the final portion of the summary. By drawing attention to this issue, our study contributes to ongoing efforts to enhance the reliability and factual consistency of LLM-generated summaries.
Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
Appendix A Details for Generating Summaries
Building upon Section 2, this section details our approach to generating long summaries and presents an analysis of the generated outputs.
Prompt
Short outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range. The specific word ranges used for prompting long-summary generation are provided in Table 5, and the prompt template employed for summary generation is shown in Table 7.
Generated Summary Length
Our experiments revealed that controlling the output length of LLMs remains a significant challenge. This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger. In particular, with longer contexts, LLMs frequently generate long summaries that contain repeated sentences or even entire paragraphs. To ensure a fair and reliable evaluation, we excluded such faulty outputs and sampled only summaries with comparable lengths for use in our analysis.
Table 6 presents the average word counts of both short and long summaries generated by each model. As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
Models Avg. Words
short long
Llama3.1-8B-Instruct 279 834
Qwen2.5-7B-Instruct 153 751
Mistral-7B-Instruct-v0.3 285 794
GPT-4o mini 172 925
LongWriter-Llama3.1-8B 398 3291
Table 6: The average word counts of both short and long summaries generated by each model on the Wikipedia 7K dataset.
Appendix B Details for Evaluation Metric
In this section, we present the experimental details of the evaluation metric used to compute sentence-level faithfulness scores in Section 3.1.
For atomic fact decomposition, we utilize the Llama3.1-8B-Instruct model with vLLM Kwon et al. (2023) [^1] https://github.com/vllm-project/vllm. For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al. (2024) [^2] https://huggingface.co/vectara/hallucination_evaluation_model.
Appendix C More Results on Varying Models
Usage of LLMs
We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace. For the GPT-4o mini model, we used gpt-4o-mini-2024-07-18. All summaries were generated using greedy decoding with float16 precision.
More Results
We report the faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset in Table 8, and those of long summaries in Table 9.
Appendix D More Results on Varying Datasets
We report comparison of faithfulness scores in short summaries generated by the Llama 3.1 model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Table 8.
We report the faithfulness scores and Sensitivity of short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 10, and those of long summaries in Table 11.
As demonstrated by the experimental results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
Appendix E Attention Weight Calculation
This section provides a more detailed and formal explanation our attention weight computation method introduced in Section 4.2.
Let the full sequence, consisting of the input context and the generated output summary, be defined as:
$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
where $L$ is the number of tokens.
We partition $x_{full}$ into $K$ non-overlapping blocks of 100 tokens each, such that each block is represented as:
$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
for $k\in\{1,2,...,K\}$.
Let the output summary contain three target sentences: the first, middle, and last sentence, denoted respectively as:
$x^{first}$ =$=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{second}$ =$=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{third}$ =$=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
We define the average attention weight from block $x^{k}$ to each of the three sentences as follows:
$\begin{gathered}\textrm{{Attn}}_{first}(k)=\frac{1}{100\cdot{T_{1}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{1}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{first})\\ \textrm{{Attn}}_{middle}(k)=\frac{1}{100\cdot{T_{2}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{2}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{middle})\\ \textrm{{Attn}}_{last}(k)=\frac{1}{100\cdot{T_{3}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{3}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{last})\end{gathered}$ (7)
where $\textrm{{attn}}(x_{i}\to x_{j})$ denotes the attention weight from token $x_{i}$ to token $x_{j}$ when generating the output.
These measures quantify the extent to which each 100-token block in the full sequence contributes to the generation of the respective sentences in the output summary.
Appendix F Details for Mitigation Methods
In this section, we provide the experimental details of the mitigation methods described in Section 5.
For the BooookScore method, the input is divided into 2048-token chunks, summaries are generated for each chunk using the Llama3.1-8B-Instruct model, and the partial summaries are then hierarchically merged into a final summary. For LongWriter, we employed the LongWriter-Llama3.1-8B model [^6] https://huggingface.co/THUDM/LongWriter-llama3.1-8b with greedy decoding and bfloat16 precision.
Input Prompt
Write an accurate and engaging summary for the given text in range of {{words_range}} words using only the provided passage (might be irrelevant). Use an unbiased and journalistic tone. Text: {{Text}}
Table 7: Prompt template used to generate summaries from the original document in Section 2.
Figure 8: Comparison of faithfulness scores in short summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.73 0.73 0.75 0.74 0.71 3.0
2K 0.77 0.76 0.75 0.73 0.73 2.25
3K 0.79 0.80 0.76 0.76 0.69 8.75
4K 0.81 0.82 0.81 0.79 0.76 4.75
5K 0.80 0.79 0.79 0.76 0.75 3.5
6K 0.81 0.80 0.80 0.76 0.71 8.25
7K 0.85 0.76 0.78 0.78 0.70 9.25
8K 0.80 0.76 0.73 0.74 0.63 12.75
Qwen2.5-7B-Instruct 1K 0.78 0.75 0.77 0.72 0.69 6.5
2K 0.83 0.82 0.74 0.72 0.74 3.75
3K 0.82 0.78 0.69 0.81 0.72 5.5
4K 0.85 0.77 0.78 0.72 0.74 4.0
5K 0.83 0.82 0.79 0.80 0.74 7.0
6K 0.79 0.75 0.77 0.75 0.70 6.5
7K 0.80 0.82 0.77 0.80 0.75 4.75
8K 0.84 0.81 0.78 0.82 0.77 4.25
GPT-4o mini 1K 0.79 0.69 0.76 0.73 0.65 9.25
2K 0.79 0.75 0.73 0.66 0.66 7.25
3K 0.84 0.81 0.80 0.73 0.71 8.5
4K 0.74 0.78 0.75 0.72 0.68 6.75
5K 0.86 0.81 0.79 0.75 0.68 12.25
6K 0.77 0.80 0.76 0.77 0.67 10.5
7K 0.87 0.85 0.75 0.77 0.68 13.0
8K 0.80 0.80 0.76 0.77 0.68 10.25
Mistral-7B-Instruct-v0.3 1K 0.83 0.77 0.77 0.75 0.80 -2.0
2K 0.74 0.80 0.80 0.73 0.67 9.75
3K 0.76 0.75 0.78 0.78 0.70 6.75
4K 0.79 0.82 0.75 0.80 0.74 5.0
5K 0.83 0.83 0.84 0.81 0.78 4.75
6K 0.82 0.81 0.82 0.84 0.79 3.25
7K 0.83 0.82 0.88 0.87 0.81 4.0
8K 0.85 0.84 0.83 0.82 0.77 6.5
Table 8: Comparison of faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.76 0.75 0.69 0.68 0.68 4.0
2K 0.73 0.76 0.73 0.70 0.63 10.0
3K 0.71 0.72 0.72 0.74 0.62 10.25
4K 0.71 0.75 0.70 0.70 0.67 4.5
5K 0.71 0.75 0.79 0.79 0.70 6.0
6K 0.69 0.76 0.73 0.76 0.61 12.5
7K 0.78 0.79 0.75 0.76 0.67 10.0
8K 0.74 0.79 0.74 0.77 0.67 9.0
Qwen2.5-7B-Instruct 1K 0.83 0.76 0.79 0.73 0.70 7.75
2K 0.81 0.79 0.78 0.78 0.70 9.0
3K 0.81 0.78 0.80 0.73 0.70 8.0
4K 0.77 0.76 0.74 0.76 0.69 6.75
5K 0.82 0.81 0.77 0.76 0.71 8.0
6K 0.81 0.78 0.78 0.75 0.67 11.0
7K 0.81 0.74 0.75 0.76 0.75 1.5
8K 0.79 0.78 0.79 0.78 0.73 5.5
GPT-4o mini 1K 0.77 0.74 0.69 0.69 0.69 3.3
2K 0.74 0.72 0.70 0.66 0.63 7.5
3K 0.75 0.70 0.69 0.65 0.58 11.8
4K 0.74 0.71 0.68 0.69 0.64 6.5
5K 0.73 0.72 0.72 0.68 0.60 11.3
6K 0.74 0.72 0.66 0.65 0.61 8.3
7K 0.73 0.70 0.69 0.66 0.62 7.5
8K 0.76 0.70 0.72 0.69 0.61 10.8
Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
2K 0.75 0.80 0.84 0.82 0.81 -0.8
3K 0.81 0.80 0.82 0.84 0.83 -1.3
4K 0.79 0.81 0.83 0.81 0.81 0.0
5K 0.80 0.84 0.86 0.85 0.79 4.8
6K 0.82 0.87 0.87 0.87 0.86 -0.3
7K 0.82 0.84 0.82 0.83 0.85 -2.3
8K 0.87 0.83 0.87 0.89 0.81 5.5
Table 9: Comparison of faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.73 0.73 0.75 0.74 0.71 3.0
2K 0.77 0.76 0.75 0.73 0.73 2.25
3K 0.79 0.80 0.76 0.76 0.69 8.75
4K 0.81 0.82 0.81 0.79 0.76 4.75
5K 0.80 0.79 0.79 0.76 0.75 3.5
6K 0.81 0.80 0.80 0.76 0.71 8.25
7K 0.85 0.76 0.78 0.78 0.70 9.25
8K 0.80 0.76 0.73 0.74 0.63 12.75
Arxiv 1K 0.69 0.76 0.79 0.81 0.71 5.25
2K 0.75 0.76 0.80 0.76 0.70 6.75
3K 0.79 0.86 0.84 0.83 0.70 13.0
4K 0.79 0.83 0.83 0.83 0.75 7.0
5K 0.82 0.84 0.84 0.83 0.79 4.25
6K 0.81 0.83 0.81 0.81 0.81 0.5
7K 0.85 0.86 0.86 0.86 0.76 9.75
8K 0.78 0.83 0.83 0.81 0.77 4.25
Pubmed 1K 0.81 0.83 0.81 0.73 0.79 0.5
2K 0.82 0.84 0.78 0.81 0.74 7.25
3K 0.83 0.84 0.81 0.83 0.76 6.75
4K 0.81 0.82 0.78 0.80 0.76 4.25
5K 0.81 0.80 0.83 0.82 0.73 8.5
6K 0.79 0.80 0.86 0.83 0.78 4.0
7K 0.80 0.81 0.82 0.79 0.78 2.5
8K 0.83 0.84 0.81 0.84 0.75 8.0
Govreport 1K 0.72 0.79 0.80 0.82 0.76 2.25
2K 0.83 0.82 0.86 0.83 0.77 6.5
3K 0.83 0.84 0.83 0.82 0.81 2.0
4K 0.82 0.76 0.83 0.78 0.78 1.75
5K 0.83 0.81 0.86 0.82 0.82 1.0
6K 0.84 0.83 0.87 0.86 0.82 3.0
7K 0.83 0.86 0.89 0.89 0.79 7.75
8K 0.84 0.82 0.87 0.85 0.81 3.5
Table 10: Comparison of faithfulness scores and Sensitivity for short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.76 0.75 0.69 0.68 0.68 4.0
2K 0.73 0.76 0.73 0.70 0.63 10.0
3K 0.71 0.72 0.72 0.74 0.62 10.25
4K 0.71 0.75 0.70 0.70 0.67 4.5
5K 0.71 0.75 0.79 0.79 0.70 6.0
6K 0.69 0.76 0.73 0.76 0.61 12.5
7K 0.78 0.79 0.75 0.76 0.67 10.0
8K 0.74 0.79 0.74 0.77 0.67 9.0
Arxiv 1K 0.77 0.82 0.81 0.80 0.74 6.0
2K 0.76 0.78 0.80 0.72 0.71 5.5
3K 0.75 0.79 0.81 0.78 0.68 10.25
4K 0.71 0.81 0.74 0.77 0.73 2.75
5K 0.77 0.83 0.85 0.81 0.74 7.5
6K 0.69 0.82 0.79 0.75 0.75 1.25
7K 0.86 0.86 0.80 0.80 0.77 6.0
8K 0.79 0.81 0.85 0.82 0.73 8.75
Pubmed 1K 0.75 0.85 0.84 0.80 0.78 3.0
2K 0.76 0.83 0.82 0.83 0.72 9.0
3K 0.81 0.85 0.86 0.78 0.76 6.5
4K 0.77 0.82 0.79 0.79 0.68 11.25
5K 0.78 0.81 0.81 0.78 0.70 9.5
6K 0.80 0.87 0.83 0.77 0.74 7.75
7K 0.85 0.85 0.84 0.79 0.77 6.25
8K 0.80 0.79 0.81 0.77 0.67 12.25
Govreport 1K 0.80 0.80 0.85 0.87 0.78 5.0
2K 0.82 0.85 0.83 0.85 0.78 5.75
3K 0.81 0.86 0.89 0.81 0.82 2.25
4K 0.76 0.86 0.86 0.82 0.77 5.5
5K 0.82 0.85 0.88 0.82 0.78 6.25
6K 0.86 0.88 0.88 0.87 0.80 7.25
7K 0.86 0.91 0.89 0.88 0.83 5.5
8K 0.84 0.87 0.89 0.84 0.80 6.0
Table 11: Comparison of faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.