### v2

# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to the source material remains a significant challenge due to the generation of hallucinations by LLMs. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucinations within generated text, particularly in long outputs. In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization
Joonho Yang1, Seunghyun Yoon2, Hwan Chang1, Byeongjeong Kim1, Hwanhee Lee1thanks: Corresponding author. 1Department of Artificial Intelligence, Chung-Ang University, 2Adobe Research, USA {plm3332, hwanchang, michael97k, hwanheelee}@cau.ac.kr, syoon@adobe.com
1 Introduction
Recent advancements in Large Language Models (LLMs) have pushed the boundaries of human-like language generation, particularly in tasks like text summarization Chang et al. (2024b). With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Song et al. (2025). This capability is transforming how we interact with large volumes of text, making information more accessible Achiam et al. (2023); Team et al. (2024).
Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
As LLMs continue to evolve, a paradigm shift is emerging from short-form to long-form generation, enabled by the ability to process extended contexts Wu et al. (2025a). Long-output generation is critical for complex reasoning tasks like long Chain-of-Thought prompting Jaech et al. (2024) and language agents Sumers et al. (2024); Wang et al. (2024), inherently requiring coherent, contextually grounded long responses. While recent benchmarks evaluate long-generation capabilities Wu et al. (2025b); Ye et al. (2025), they often lack contextual grounding, overlooking a crucial aspect of long-form generation: context-aware faithfulness.
A persistent and critical challenge faced by LLMs is the phenomenon of hallucination, wherein the generated content is unfaithful to or unsupported by the input context Huang et al. (2025). While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the positional distribution of factual errors within the generated sequence. Related research has explored positional biases in how models process input contexts, notably the Lost in the Middle phenomenon Liu et al. (2024b); Wan et al. (2025). However, understanding the distribution of errors within the generated output itself, particularly for long responses from long contexts, is equally crucial for effective diagnosis and mitigation.
Motivated by this critical research gap, our work provides the first dedicated investigation into the positional distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization. Unlike prior work primarily addressing long input contexts or short outputs Zhang et al. (2024a); Bishop et al. (2024), we tackle long context-aware long response generation, which requires processing lengthy inputs and maintaining faithfulness over a significantly long output sequence. Specifically, we pose a core research question: How frequently do hallucinations occur in long-form generation tasks such as document summarization, and when they do, where in the output are they most likely to appear?
Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness significantly decreases towards the end for long summaries (e.g., 800 words). Contrary to expectations of uniform distribution, this distinct positional bias highlights a critical vulnerability in LLMs when generating extended text. To systematically analyze this phenomenon and its implications, our research addresses three core questions:
  1. RQ1.
    Where Do Hallucinations Most Frequently Occur? (§3)
  2. RQ2.
    What Factors Contribute to Hallucination Concentration at the Last Part? (§4)
  3. RQ3.
    How Can We Mitigate the Hallucination in the Last Part? (§5)
To answer these questions, we empirically characterize the positional distribution of hallucinations (RQ1), investigate their generative causes (RQ2), and propose mitigation strategies (RQ3). Our results underscore the importance of the positional dimension of errors in building robust long-form generation systems.
Our contributions can be summarized as follows:
  • We provide the first empirical characterization of the Hallucinate at the Last phenomenon in LLM-based long response generation, particularly in long document summarization.
  • We offer initial insights and analysis into the potential factors contributing to this late-stage hallucination.
  • We investigate mitigation strategies specifically tailored to address positional hallucination.
2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents, varying both the context length and the output length. For all experiments, we generate summaries using greedy decoding. The prompt is provided in Appendix A.
2.1 Input Context Lengths
To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 8K tokens. For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
2.2 Output Length Categories
We define two categories for output length to study the Hallucinate at the Last phenomenon:
  • Standard Summary: Summaries with a length between 100 and 200 words.  [^1] The length of the standard summaries in our study was determined based on the average length of the reference summaries in the dataset, which is employed in Section 4.
  • Long Summary: Summaries with a length up to 30% of the input context length.
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both standard and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix A for more details.
2.3 Overall Faithfulness
As a preliminary step, we measure the overall faithfulness of the generated long summaries with FineSurE Song et al. (2024) to ensure consistent quality across different context lengths, prior to our main positional analysis. Table 1 shows results for Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset (1K–8K context lengths), evaluated using FineSurE with GPT-4 Achiam et al. (2023). FineSurE assesses each sentence against the source, identifying error types when inconsistencies occur. The results align with those on CNNDM Hermann et al. (2015), where Llama3-70B-Instruct achieved 85.5% faithfulness. We next conduct a finer-grained positional analysis in Section 3 for various setups.
Context Length Faithfulness(%)
1K 83.3
2K 86.7
3K 78.9
4K 82.9
5K 86.6
6K 85.9
7K 83.9
8K 86.8
Table 1: Faithfulness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
3 Where Do Hallucinations Most Frequently Occur?
In this section, we empirically investigate where hallucinations most frequently occur in long document summarization. We analyze the positional distribution of factual errors by examining faithfulness across varying models, datasets, and output positions.
Figure 2: Comparison of faithfulness scores for standard summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
3.1 Experimental Setup
Models
We evaluate our approach on multiple state-of-the-art LLMs, including Llama3.1-8B-Instruct, Gemma3-12B-Instruct Team et al. (2025), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini, to ensure model-agnostic findings. For Gemma, we retain the default sliding window attention context length of 1024 tokens, while for Qwen, we modify the default setting of over 131K tokens to 16K tokens.
Datasets
To examine whether the phenomenon is domain-specific, we analyze the positional distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation.
3.2 Evaluation Metric
Faithfulness
To enable an in-depth analysis of positional bias within each generated summary, we use the MiniCheck-Flan-T5-Large Tang et al. (2024) model to evaluate faithfulness. MiniCheck is particularly suitable for this granular task, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al. (2023); Jing et al. (2024); Yu et al. (2024).
Let $s_{i}$ be the $i$-th sentence in the generated summary $S$. Each sentence in the generated summary is decomposed into a set of atomic facts $A_{i}$ by utilizing LLM, followed by a filtering process to remove unnecessary atomic facts Yang et al. (2024b), which makes:
$A_{i}=\{a_{i1},a_{i2},...,a_{iN_{i}}\}$ (1)
Using a MiniCheck model, which is denoted as $\mathcal{M}$, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner. Each atomic fact is assigned the highest score among its comparisons:
$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\mathcal{M}(d_{m},a_{ij})$ (2)
Finally, we use the average of the scores for a set of filtered atomic facts in a sentence as its faithfulness score:
$\textrm{Faithfulness}(s_{i})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\textrm{score}(a_{ij})$ (3)
Additional experimental details and an example of the evaluation procedure are provided in Appendix B.
Furthermore, we verify the quality of atomic facts generated by the LLM via human evaluation in Table 2 for atomic facts extracted from long summaries generated by Llama.
Context Length Validation(%)
1K 98.55
2K 99.08
3K 97.62
4K 95.92
5K 99.01
6K 96.18
7K 99.12
8K 99.64
Table 2: Human evaluation of the quality (completeness) of atomic facts. For each context length, we randomly sampled 10 summaries and conducted the evaluation manually. Each atomic fact was annotated as either true or false depending on whether it is present in the summary.
Sensitivity
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Specifically, following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins or segments, and the average faithfulness score of the atomic facts within each bin is computed. We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins. A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise. The larger the sensitivity, the more pronounced the hallucination at the end of the output.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct Overall Avg. 0.81 0.82 0.79 0.80 0.75 5.5
12K 0.80 0.79 0.66 0.63 0.69 9.5
16K 0.86 0.83 0.85 0.83 0.79 5.25
Llama3.1-70B-Instruct Overall Avg. 0.83 0.82 0.83 0.81 0.77 5.25
Gemma3-4B-Instruct Overall Avg. 0.85 0.82 0.83 0.85 0.79 4.75
Gemma3-12B-Instruct Overall Avg. 0.85 0.83 0.84 0.83 0.79 4.75
GPT-4o mini Overall Avg. 0.83 0.81 0.81 0.78 0.76 4.75
Qwen2.5-7B-Instruct Overall Avg. 0.86 0.83 0.83 0.82 0.84 -0.5
Table 3: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset. Overall Avg. denotes the average faithfulness score aggregated across context lengths from 1K to 8K.
3.3 Hallucinate at the Last
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.92 0.90 0.91 0.87 0.68 22.00
2K 0.82 0.74 0.83 0.72 0.71 6.75
3K 0.94 0.94 0.81 0.86 0.68 20.75
4K 0.92 0.92 0.88 0.87 0.82 7.75
5K 0.91 0.88 0.89 0.78 0.70 16.50
6K 0.86 0.84 0.79 0.89 0.73 11.50
7K 0.97 0.89 0.82 0.94 0.66 24.50
8K 0.93 0.90 0.86 0.79 0.65 22.00
Table 4: Human evaluation of faithfulness scores and sensitivity for long summaries generated by Llama on the Wikipedia dataset. Each atomic fact was annotated as either true or false depending on whether it is present in the document. The results were manually evaluated and aggregated.
The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across various contexts and summary lengths. The decline becomes most pronounced when the context length reaches 8K and the summaries are long.
We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively. As shown in Figure 2, standard summaries generated by the three models, excluding Qwen, consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end. Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length. Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries except for Qwen, as shown in Figure 3. For these models, as the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below a faithfulness score of 0.70 when 45 sentences are generated from an 8K context. Qwen, in contrast, often exhibits the largest drop in faithfulness in the middle of the summary, after which the scores recover toward the end.
Tables 3 report the faithfulness and sensitivity of various models. We aggregate the faithfulness scores across context lengths from 1K to 8K and report the overall average. All models except Qwen show markedly low faithfulness scores in the final bin, reflecting high sensitivity. Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size or context length. Qwen, however, is the only model that records sensitivity values below zero. Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity, and we provide a detailed analysis of these contrasting patterns in Section 4. Supplementary results across context lengths and methods are provided in Appendix C.
The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, with a marked decline once the context length exceeds 7K.
We present the positional faithfulness discrepancy of long summaries generated by the Llama model across different datasets in Figure 4. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores tend to decrease toward the final segments as the output length increases. In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline, with the deterioration becoming especially pronounced once the context length exceeds 7K. Additional sensitivity results across diverse datasets are provided in Appendix D.
Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
4 Why Do Hallucinations Frequently Occur at the Last?
This section investigates the underlying factors that contribute to the Hallucinate at the Last phenomenon, exploring two main hypotheses. The first attributes the phenomenon to the inherent nature of summarization: key information is typically concentrated at the beginning of the summary, with less important content appearing toward the end. The second posits that as LLMs generate longer outputs, they increasingly attend to previously generated tokens rather than the original input context, resulting in a shift in attention distribution.
4.1 Is it Intrinsic to Summarization?
This hypothesis draws on the intrinsic structure of both human- and model-generated summaries. In typical summarization domains such as news, scientific articles, and Wikipedia-style documents, key information is predominantly introduced early in the summary Hermann et al. (2015); Narayan et al. (2018); Cohan et al. (2018), reflecting the lead bias observed in source texts, where the opening contains the most salient content Kim et al. (2019); Zhao et al. (2022); Ravaut et al. (2024). Building on this, our study investigates whether this structural bias is also reflected in decreasing factual consistency within generated outputs by comparing positional faithfulness scores between human-written and model-generated summaries.
Figure 5: Faithfulness scores of reference summaries and those generated by Llama on the CNNDM dataset.
Experimental Setup
We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens. We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
Results & Analysis
Figure 5 compares positional faithfulness between human-written and model-generated CNNDM summaries. We observe that the faithfulness of the generated summaries declines steadily, falling below 0.6 in the final segment, whereas reference summaries dip around the middle but recover. This suggests that Hallucination at the Last is not solely due to summarization’s inherent structure.
4.2 Is it Intrinsic to Attention?
Prior research on LLMs has demonstrated that attention weight distributions are closely correlated with the generation process, influencing output coherence Dong et al. (2021); Zhang et al. (2024b). Recent studies on Large Vision-Language Models (LVLMs) further suggest that hallucinations frequently occur in the later parts of generated text Liu et al. (2024a); Lee et al. (2024); Min et al. (2025). This phenomenon has been attributed to a shift in attention: as text generation progresses, attention weights increasingly favor previously generated text tokens over image tokens. Inspired by this observation, we investigate whether a similar trend exists in LLMs by analyzing how attention weights on generated tokens evolve as output length increases.
Experimental Setup
Unlike previous studies that focus on attention to the first generated token Hsieh et al. (2024), we compute attention weights at the sentence level. Moreover, we segment the full sequence into chunks of 100 tokens. Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence. Details of the attention weight calculation are provided in Appendix E.
Results & Analysis
Figure 6 presents a visualization of average attention weights across sentences. As shown in the results, the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 3, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Qwen model assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination.
This consistent distribution in Qwen can be attributed to its sliding window attention mechanism Child et al. (2019), which maintains a balanced focus across the context and mitigates the downstream attention concentration that triggers hallucinations. Table 5 compares the Qwen model with and without sliding window attention. In this setting, the sliding window context length is fixed at 16000 tokens. The results clearly show that sliding window attention omission increases sensitivity across all context and summary lengths, up to 9.75. They demonstrate that employing sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
(a) Average attention weights of summaries generated by Llama.
(b) Average attention weights of summaries generated by Qwen.
Figure 6: Average attention weights of two models exhibiting contrasting trends in Figure 3. The dashed line separates the input context from the generated output.
5 How to Mitigate the Hallucinate at the Last?
In the preceding section, we demonstrated that the Hallucinate at the Last phenomenon could be partially mitigated through the use of sliding window attention. In the following analysis, we investigate strategies to address this issue in models that are not trained with sliding window attention. To this end, we apply four methods that may mitigate the phenomenon.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Standard Summary
Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
2K 0.90 0.89 0.83 0.78 0.84 1.0
3K 0.87 0.84 0.79 0.83 0.82 1.25
4K 0.91 0.88 0.90 0.79 0.89 -2.0
5K 0.86 0.89 0.85 0.76 0.89 -5.0
6K 0.93 0.91 0.92 0.80 0.90 -1.0
7K 0.90 0.89 0.89 0.84 0.82 6.0
8K 0.89 0.93 0.88 0.89 0.92 -2.25
Without 1K 0.88 0.79 0.74 0.76 0.84 -4.75(+1.5)
2K 0.90 0.89 0.83 0.78 0.84 1.0(+0.0)
Sliding 3K 0.88 0.86 0.82 0.85 0.83 2.25(+1.0)
4K 0.91 0.88 0.87 0.83 0.83 4.25(+6.25)
Winodw 5K 0.87 0.87 0.81 0.82 0.80 4.25(+9.25)
6K 0.93 0.90 0.91 0.85 0.86 3.75(+4.75)
Attention 7K 0.91 0.91 0.89 0.84 0.79 9.75(+3.75)
8K 0.88 0.93 0.89 0.89 0.92 -2.25(+0.0)
Long Summary
Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
2K 0.83 0.89 0.84 0.83 0.82 2.75
3K 0.83 0.84 0.86 0.79 0.84 -1.0
4K 0.87 0.83 0.79 0.83 0.84 -1.0
5K 0.87 0.89 0.89 0.80 0.85 1.25
6K 0.89 0.81 0.78 0.78 0.82 -0.5
7K 0.86 0.78 0.79 0.83 0.84 -2.5
8K 0.86 0.85 0.85 0.85 0.85 0.25
Without 1K 0.89 0.77 0.83 0.84 0.80 3.25(+4.25)
2K 0.83 0.89 0.83 0.80 0.78 5.75(+3.0)
Sliding 3K 0.83 0.84 0.86 0.79 0.84 -1.0(+0.0)
4K 0.87 0.83 0.79 0.83 0.84 -1.0(+0.0)
Winodw 5K 0.86 0.90 0.83 0.79 0.80 4.5(+3.25)
6K 0.89 0.82 0.84 0.83 0.78 6.5(+7.0)
Attention 7K 0.86 0.80 0.78 0.83 0.82 -0.25(+2.25)
8K 0.88 0.85 0.85 0.86 0.82 4.0(+3.75)
Table 5: Faithfulness scores and sensitivity for standard and long summaries generated by Qwen2.5-7B-Instruct on the Wikipedia dataset, with and without sliding window attention. Note that higher sensitivity indicates stronger hallucination effects toward the end of the output.
Experimental Setup
In this experiment, we generate Wikipedia summaries with a 7K context using four methods, taking Llama3.1-8B-Instruct as the baseline. For comparison, we evaluate the following four methods: (See Appendix F for more experimental details.)
  • BooookScore Chang et al. (2024a) segments the input context into chunks, generates summaries for each chunk individually, and then merges the partial summaries.
  • MInference Jiang et al. (2024) employs sparse attention mechanisms to efficiently process long input sequences.
  • LongWriter-Llama3.1-8B Bai et al. (2025) is a model fine-tuned on a long-output dataset and further enhanced using DPO Rafailov et al. (2023).
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024).
Methods Generated Summary Bins Sensitivity
1 2 3 4 5
Llama3.1-8B 0.74 0.75 0.75 0.73 0.64 10.3
+ BooookScore 0.73 0.77 0.78 0.80 0.75 2.0
+ MInference 0.77 0.78 0.78 0.71 0.69 7.0
+ LongWriter 0.85 0.77 0.81 0.63 0.70 6.5
+ AdaCAD 0.79 0.82 0.77 0.81 0.64 15.8
Table 6: Faithfulness scores and sensitivity for long summaries generated by different mitigation methods on the Wikipedia dataset.
Results & Analysis
We report the faithfulness scores across output bins and the sensitivity results in Table 6, and the faithfulness scores across increasing output lengths for each method in Figure 7. As shown in Table 6, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. Nevertheless, while BooookScore serves as a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation. This highlights the need for future work to develop methods that maintain faithfulness when generating long text directly from long contexts.
6 Related Work
The problem of hallucination in LLMs has been a significant area of research. Prior work has largely focused on developing methods for detecting factual inconsistencies in generated text Chuang et al. (2024); Hu et al. (2024); Kim et al. (2024); Zhong and Litman (2025) and proposing strategies to mitigate their occurrence, often through improved training Zhang et al. (2023); Wan et al. (2023), decoding Shi et al. (2024); Wang et al. (2025), or prompting Zhou et al. (2023). While effective in reducing overall hallucination rates, these studies typically do not analyze the within-sequence distribution of errors.
The concept of positional bias in generation errors has been observed in related domains. Notably, studies on Large Vision-Language Models (LVLMs) have indicated that hallucinations tend to increase towards the end of generated textual descriptions of images Liu et al. (2024a); Lee et al. (2024); Min et al. (2025).
Recently, several studies have begun to explore long-output generation in LLMs Wu et al. (2025b); Ye et al. (2025). However, these works focus on procedural generation tasks that are not grounded in contextual input, limiting their applicability to real-world summarization scenarios. A notable study on multi-document summarization aligns with our findings Belém et al. (2025); Zhao et al. (2025). Nevertheless, it does not address positional context-aware faithfulness over extended summary lengths.
Our work presents the first dedicated study of where hallucinations occur in LLM-based long document summarization, moving beyond detection to uncover positional patterns, underlying causes, and mitigation strategies.
Figure 7: Faithfulness scores for long summaries generated by different mitigation methods on the Wikipedia dataset across increasing output lengths.
7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into positionally-aware generation techniques.
Limitations
In this study, we explored four domains, primarily because the corresponding datasets provided input contexts of the desired length. However, we were not able to investigate domains such as books, dialogues, movie scripts, or meeting transcripts. Books were excluded due to their excessive length, while the other domains lacked datasets with suitable input lengths or sufficient sample sizes.
Despite these limitations, we believe this work offers an important foundation for future research in the relatively underexplored domain of long-form output generation, particularly in summarization.
Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
Appendix A Details for Generating Summaries
Building upon Section 2, this section details our approach to generating long summaries and presents an analysis of the generated outputs.
Prompt
Standard outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range. The specific word ranges used for prompting long-summary generation are provided in Table 7, and the prompt template employed for summary generation is shown in Table 10.
Context Length Words Range
1K 200 to 250
2K 400 to 500
3K 600 to 750
4K 800 to 1000
5K 1000 to 1250
6K 1200 to 1500
7K 1400 to 1750
8K 1600 to 2000
Table 7: Words range used in the prompt for long output generation.
Generated Summary Length
Our experiments revealed that controlling the output length of LLMs remains a significant challenge. This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger. In particular, with longer contexts, LLMs frequently generate long summaries that contain repeated sentences or even entire paragraphs. To ensure a fair and reliable evaluation, we excluded such faulty outputs and sampled only summaries with comparable lengths for use in our analysis.
Table 8 presents the average word counts of both standard and long summaries generated by each model. As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
Models Avg. Words
standard long
[…]
Table 8: The average word counts of both standard and long summaries generated by each model on the Wikipedia 7K dataset.
Appendix B Details for Model Usage
In this section, we present the experimental details for the model usage in Section 3.
Details for Evaluation Method
For atomic fact decomposition, we utilize the Llama3.1-8B-Instruct model with vLLM Kwon et al. (2023) [^2] https://github.com/vllm-project/vllm. To further illustrate the evaluation process, we present in Table 9 a concrete example that demonstrates how a given sentence is decomposed into its constituent atomic facts and subsequently assessed for faithfulness. This example clarifies the step-by-step procedure and provides transparency regarding the alignment between the generated atomic facts and the underlying reference information.
First Sentence
Sara Sorribes Tormo is a talented Spanish professional tennis player born on October 8, 1996, in Castellón de la Plana, Spain.
Atomic Facts Score
Sara Sorribes Tormo is a Spanish professional tennis player. 0.92
Sara Sorribes Tormo was born in Spain. 0.90
Sara Sorribes Tormo was born on October 8. 0.87
Sara Sorribes Tormo was born in 1996. 0.96
Sara Sorribes Tormo was born in Castellón de la Plana. 0.93
Avg. Score 0.92
Last Sentence
As Sara navigates her career, her unique blend of skill, perseverance, and spirit positions her as a promising athlete, inspiring both current and aspiring players.
Atomic Facts Score
Sara navigates her career. 0.87
Sara’s career is that of an athlete. 0.91
Sara has a unique blend of skill. 0.14
Sara has perseverance. 0.81
Sara is a promising athlete. 0.30
Sara inspires current players. 0.11
Sara inspires aspiring players. 0.15
Avg. Score 0.47
Table 9: Example of the evaluation procedure with the MiniCheck-Flan-T5-Large Model.
Details for LLMs Usage
We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Gemma3-12B-Instruct [^4] https://huggingface.co/google/gemma-3-12b-it, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace. For the GPT-4o mini model, we used gpt-4o-mini-2024-07-18. All summaries were generated using greedy decoding with float16 precision.
Appendix C More Results with Different Methods
We adopt a diverse set of methods to validate the robustness of our faithfulness evaluation framework. In this section, we present experimental results for faithfulness assessment models beyond MiniCheck, incorporating approaches based on NLI, ROUGE-L, and LLMs. Specifically, we substitute the scoring model $\mathcal{M}$ in Equation 2 of Section 3 with these alternative methods.
Minicheck Based
We utilize MiniCheck-Flan-T5-Large Model [^6] https://huggingface.co/lytang/MiniCheck-Flan-T5-Large for the main evaluation method. We report the faithfulness scores and sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 11, and those of long summaries in Table 12.
NLI Based
For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al. (2024) [^7] https://huggingface.co/vectara/hallucination_evaluation_model.
We report the faithfulness scores and sensitivity assessed by NLI model for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 13, and those of long summaries in Table 14. As observed in the experimental results, consistent with the main results, the Llama and GPT models consistently exhibit high sensitivity, whereas the Qwen model occasionally demonstrates sensitivity values even below zero.
ROUGE Based
We report the faithfulness scores and sensitivity assessed by ROUGE-L Lin (2004) for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 17, and those of long summaries in Table 18.
LLM Based
We present the error rates assessed by FineSurE framework for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 15, and those of long summaries in Table 16.
As observed in the experimental results, consistent with the main results, the Llama and GPT models consistently exhibit the highest error rates in the final bin. In contrast, the Qwen model does not yield the highest error rates in the final bin for both the cases of standard and long summaries.
Appendix D More Results on Varying Datasets
We report comparison of faithfulness scores in standard summaries generated by the Llama model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Figure 8.
Minicheck Based
We report the faithfulness scores and sensitivity assessed by MiniCheck-Flan-T5-Large model of standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 19, and those of long summaries in Table 20.
The experimental results reveal that in the case of standard summaries, sensitivities below zero are frequently observed; notably, in the Pubmed dataset, more than half of the values fall below zero. In contrast, such negative sensitivities are rarely observed in long summaries, with the Pubmed dataset in particular showing no sensitivities below zero, unlike the standard summary setting.
NLI Based
We report the faithfulness scores and sensitivity assessed by NLI model for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 21, and those of long summaries in Table 22.
As demonstrated by the experimental results, consistent with the main results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
Appendix E Attention Weight Calculation
This section provides a more detailed and formal explanation of our attention weight computation method introduced in Section 4.
Let the full sequence, consisting of the input context and the generated output summary, be defined as:
$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
where $L$ is the number of tokens.
We partition $x_{full}$ into $K$ non-overlapping blocks of 100 tokens each, such that each block is represented as:
$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
for $k\in\{1,2,...,K\}$.
Let the output summary contain three target sentences: the first, middle, and last sentence, denoted respectively as:
$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{second}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{third}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
We define the average attention weight from block $x^{k}$ to each of the three sentences as follows:
$\begin{gathered}\textrm{{Attn}}_{first}(k)=\frac{1}{100\cdot{T_{1}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{1}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{first})\\ \textrm{{Attn}}_{middle}(k)=\frac{1}{100\cdot{T_{2}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{2}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{middle})\\ \textrm{{Attn}}_{last}(k)=\frac{1}{100\cdot{T_{3}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{3}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{last})\end{gathered}$ (7)
where $\textrm{{attn}}(x_{i}\to x_{j})$ denotes the attention weight from token $x_{i}$ to token $x_{j}$ when generating the output.
These measures quantify the extent to which each 100-token block in the full sequence contributes to the generation of the respective sentences in the output summary.
Appendix F Details for Mitigation Methods
In this section, we provide the experimental details of the mitigation methods described in Section 5.
For the BooookScore method, the input is divided into 2048-token chunks, summaries are generated for each chunk using the Llama3.1-8B-Instruct model, and the partial summaries are then hierarchically merged into a final summary. For LongWriter, we employed the LongWriter-Llama3.1-8B model [^8] https://huggingface.co/THUDM/LongWriter-llama3.1-8b with greedy decoding and bfloat16 precision.
Appendix G Use of Large Language Model
This paper primarily utilized the AI assistant (ChatGPT-5) for grammar correction and refinement of clarity. It was not employed for content generation or research ideation.
Input Prompt
Write an accurate and engaging summary for the given text in range of {{words_range}} words using only the provided passage (might be irrelevant). Use an unbiased and journalistic tone. Text: {{Text}}
Table 10: Prompt template used to generate summaries from the original document in Section 2.
Figure 8: Comparison of faithfulness scores in standard summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.81 0.79 0.85 0.82 0.79 2.75
2K 0.84 0.79 0.81 0.81 0.77 4.25
3K 0.87 0.85 0.84 0.83 0.79 5.75
4K 0.91 0.90 0.88 0.86 0.88 0.75
5K 0.85 0.85 0.86 0.83 0.83 1.75
6K 0.89 0.85 0.85 0.83 0.81 4.5
7K 0.91 0.81 0.85 0.87 0.81 5.0
8K 0.87 0.84 0.82 0.82 0.79 4.75
Llama3.1-70B-Instruct 1K 0.87 0.83 0.86 0.80 0.77 7.0
2K 0.92 0.81 0.83 0.83 0.72 12.75
3K 0.90 0.82 0.86 0.77 0.83 0.75
4K 0.94 0.85 0.85 0.86 0.77 10.5
5K 0.91 0.86 0.85 0.83 0.89 -2.75
6K 0.89 0.85 0.84 0.87 0.83 3.25
7K 0.87 0.88 0.90 0.83 0.84 3.0
8K 0.92 0.91 0.84 0.84 0.85 2.75
Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
2K 0.90 0.89 0.83 0.78 0.84 1.0
3K 0.87 0.84 0.79 0.83 0.82 1.25
4K 0.91 0.88 0.90 0.79 0.89 -2.0
5K 0.86 0.89 0.85 0.76 0.89 -5.0
6K 0.93 0.91 0.92 0.80 0.90 -1.0
7K 0.90 0.89 0.89 0.84 0.82 6.0
8K 0.89 0.93 0.88 0.89 0.92 -2.25
GPT-4o mini 1K 0.85 0.78 0.81 0.80 0.77 4.0
2K 0.87 0.86 0.87 0.82 0.80 5.5
3K 0.89 0.88 0.86 0.85 0.80 7.0
4K 0.79 0.80 0.86 0.83 0.82 0.0
5K 0.91 0.86 0.88 0.87 0.82 6.0
6K 0.85 0.89 0.85 0.90 0.85 2.25
7K 0.92 0.87 0.84 0.90 0.84 4.25
8K 0.90 0.90 0.84 0.85 0.82 5.25
Gemma3-4B-Instruct 1K 0.88 0.89 0.85 0.71 0.77 6.25
2K 0.81 0.91 0.89 0.84 0.83 3.25
3K 0.80 0.88 0.87 0.92 0.77 9.75
4K 0.85 0.87 0.92 0.88 0.77 11.0
5K 0.90 0.83 0.83 0.92 0.82 5.0
6K 0.93 0.92 0.89 0.93 0.80 11.75
7K 0.91 0.91 0.90 0.83 0.81 7.75
8K 0.89 0.88 0.88 0.91 0.84 5.0
Gemma3-12B-Instruct 1K 0.83 0.84 0.81 0.77 0.78 3.25
2K 0.83 0.74 0.79 0.85 0.72 8.25
3K 0.85 0.82 0.77 0.85 0.76 6.25
4K 0.91 0.87 0.93 0.85 0.84 5.0
5K 0.85 0.87 0.93 0.83 0.81 6.0
6K 0.94 0.89 0.93 0.90 0.87 4.5
7K 0.91 0.85 0.86 0.82 0.77 9.0
8K 0.90 0.85 0.80 0.85 0.78 7.0
Table 11: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.78 0.82 0.75 0.75 0.75 2.5
2K 0.82 0.79 0.75 0.76 0.73 5.0
3K 0.77 0.81 0.80 0.81 0.71 8.75
4K 0.79 0.80 0.73 0.77 0.74 3.25
5K 0.80 0.83 0.83 0.84 0.79 3.5
6K 0.80 0.82 0.77 0.83 0.72 8.5
7K 0.84 0.87 0.84 0.79 0.77 6.5
8K 0.84 0.85 0.82 0.82 0.77 6.25
Llama3.1-70B-Instruct 1K 0.84 0.81 0.86 0.77 0.75 7.0
2K 0.79 0.83 0.89 0.84 0.74 9.75
3K 0.80 0.85 0.83 0.83 0.79 3.75
4K 0.83 0.81 0.85 0.82 0.74 8.75
5K 0.84 0.83 0.80 0.77 0.80 1.0
6K 0.83 0.81 0.80 0.79 0.79 1.75
7K 0.81 0.75 0.76 0.84 0.77 2.0
8K 0.87 0.86 0.84 0.85 0.79 6.5
Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
2K 0.83 0.89 0.84 0.83 0.82 2.75
3K 0.83 0.84 0.86 0.79 0.84 -1.0
4K 0.87 0.83 0.79 0.83 0.84 -1.0
5K 0.87 0.89 0.89 0.80 0.85 1.25
6K 0.89 0.81 0.78 0.78 0.82 -0.5
7K 0.86 0.78 0.79 0.83 0.84 -2.5
8K 0.86 0.85 0.85 0.85 0.85 0.25
GPT-4o mini 1K 0.85 0.81 0.81 0.74 0.80 0.25
2K 0.81 0.80 0.80 0.77 0.75 4.5
3K 0.83 0.81 0.82 0.74 0.73 7.0
4K 0.83 0.82 0.79 0.78 0.78 2.5
5K 0.84 0.82 0.84 0.81 0.75 7.75
6K 0.85 0.82 0.81 0.79 0.77 4.75
7K 0.82 0.79 0.81 0.79 0.76 4.25
8K 0.83 0.83 0.82 0.83 0.78 4.75
Gemma3-4B-Instruct 1K 0.77 0.82 0.81 0.84 0.73 8.0
2K 0.86 0.72 0.80 0.85 0.76 4.75
3K 0.87 0.82 0.84 0.84 0.79 5.25
4K 0.86 0.85 0.78 0.86 0.80 3.75
5K 0.88 0.83 0.83 0.84 0.78 6.5
6K 0.83 0.85 0.86 0.85 0.81 3.75
7K 0.82 0.82 0.88 0.86 0.82 2.5
8K 0.87 0.89 0.84 0.83 0.84 1.75
Gemma3-12B-Instruct 1K 0.90 0.83 0.80 0.82 0.81 2.75
2K 0.82 0.83 0.87 0.88 0.80 5.0
3K 0.87 0.84 0.84 0.84 0.76 8.75
4K 0.80 0.76 0.83 0.79 0.76 3.5
5K 0.86 0.79 0.86 0.79 0.80 2.5
6K 0.85 0.87 0.81 0.81 0.81 2.5
7K 0.88 0.88 0.89 0.86 0.82 5.75
8K 0.85 0.84 0.85 0.84 0.77 7.5
Table 12: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.73 0.73 0.75 0.74 0.71 3.0
2K 0.77 0.76 0.75 0.73 0.73 2.25
3K 0.79 0.80 0.76 0.76 0.69 8.75
4K 0.81 0.82 0.81 0.79 0.76 4.75
5K 0.80 0.79 0.79 0.76 0.75 3.5
6K 0.81 0.80 0.80 0.76 0.71 8.25
7K 0.85 0.76 0.78 0.78 0.70 9.25
8K 0.80 0.76 0.73 0.74 0.63 12.75
Qwen2.5-7B-Instruct 1K 0.84 0.75 0.77 0.76 0.71 7.0
2K 0.81 0.82 0.78 0.69 0.71 6.5
3K 0.78 0.76 0.72 0.73 0.68 6.75
4K 0.90 0.83 0.76 0.80 0.79 3.25
5K 0.83 0.86 0.74 0.76 0.79 0.75
6K 0.87 0.82 0.77 0.77 0.83 -2.25
7K 0.86 0.82 0.79 0.79 0.75 6.5
8K 0.85 0.88 0.84 0.80 0.76 8.25
GPT-4o mini 1K 0.79 0.69 0.76 0.73 0.65 9.25
2K 0.79 0.75 0.73 0.66 0.66 7.25
3K 0.84 0.81 0.80 0.73 0.71 8.5
4K 0.74 0.78 0.75 0.72 0.68 6.75
5K 0.86 0.81 0.79 0.75 0.68 12.25
6K 0.77 0.80 0.76 0.77 0.67 10.5
7K 0.87 0.85 0.75 0.77 0.68 13.0
8K 0.80 0.80 0.76 0.77 0.68 10.25
Gemma3-12B-Instruct 1K 0.75 0.73 0.76 0.75 0.71 3.75
2K 0.80 0.65 0.70 0.68 0.65 5.75
3K 0.78 0.76 0.70 0.75 0.66 8.75
4K 0.82 0.70 0.76 0.76 0.69 7.0
5K 0.76 0.80 0.86 0.60 0.72 3.5
6K 0.81 0.79 0.84 0.78 0.69 11.5
7K 0.83 0.81 0.78 0.79 0.74 6.25
8K 0.83 0.81 0.78 0.79 0.74 6.25
Table 13: Comparison of NLI model based faithfulness scores and Sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.76 0.75 0.69 0.68 0.68 4.0
2K 0.73 0.76 0.73 0.70 0.63 10.0
3K 0.71 0.72 0.72 0.74 0.62 10.25
4K 0.71 0.75 0.70 0.70 0.67 4.5
5K 0.71 0.75 0.79 0.79 0.70 6.0
6K 0.69 0.76 0.73 0.76 0.61 12.5
7K 0.78 0.79 0.75 0.76 0.67 10.0
8K 0.74 0.79 0.74 0.77 0.67 9.0
Qwen2.5-7B-Instruct 1K 0.83 0.74 0.76 0.75 0.68 9.0
2K 0.76 0.80 0.79 0.76 0.71 6.75
3K 0.79 0.79 0.81 0.71 0.75 2.5
4K 0.77 0.75 0.79 0.80 0.73 4.75
5K 0.83 0.84 0.79 0.76 0.76 4.5
6K 0.82 0.74 0.76 0.73 0.70 6.25
7K 0.80 0.74 0.72 0.78 0.78 -2.0
8K 0.79 0.79 0.80 0.79 0.75 4.25
GPT-4o mini 1K 0.77 0.74 0.69 0.69 0.69 3.3
2K 0.74 0.72 0.70 0.66 0.63 7.5
3K 0.75 0.70 0.69 0.65 0.58 11.8
4K 0.74 0.71 0.68 0.69 0.64 6.5
5K 0.73 0.72 0.72 0.68 0.60 11.3
6K 0.74 0.72 0.66 0.65 0.61 8.3
7K 0.73 0.70 0.69 0.66 0.62 7.5
8K 0.76 0.70 0.72 0.69 0.61 10.8
Gemma3-12B-Instruct 1K 0.81 0.79 0.72 0.72 0.73 3.0
2K 0.77 0.75 0.79 0.78 0.74 3.25
3K 0.75 0.77 0.71 0.73 0.68 6.0
4K 0.72 0.67 0.70 0.74 0.65 5.75
5K 0.77 0.78 0.80 0.75 0.71 6.5
6K 0.77 0.78 0.77 0.74 0.70 6.5
7K 0.82 0.81 0.82 0.79 0.71 10.0
8K 0.74 0.74 0.77 0.75 0.63 12.0
Table 14: Comparison of NLI model based faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.24 0.29 0.14 0.24 0.33
2K 0.21 0.21 0.16 0.24 0.37
3K 0.20 0.30 0.20 0.10 0.57
4K 0.06 0.17 0.23 0.17 0.26
5K 0.31 0.20 0.20 0.36 0.33
6K 0.21 0.25 0.29 0.14 0.50
7K 0.24 0.17 0.14 0.31 0.66
8K 0.28 0.24 0.31 0.24 0.69
Qwen2.5-7B-Instruct 1K 0.07 0.00 0.13 0.00 0.33
2K 0.13 0.20 0.13 0.00 0.13
3K 0.27 0.20 0.27 0.20 0.00
4K 0.27 0.00 0.20 0.13 0.20
5K 0.13 0.07 0.33 0.20 0.27
6K 0.07 0.07 0.27 0.13 0.20
7K 0.31 0.00 0.08 0.00 0.00
8K 0.31 0.23 0.23 0.23 0.23
GPT-4o mini 1K 0.05 0.00 0.00 0.00 0.05
2K 0.00 0.00 0.10 0.00 0.10
3K 0.05 0.00 0.00 0.00 0.10
4K 0.00 0.00 0.05 0.05 0.15
5K 0.15 0.05 0.05 0.05 0.10
6K 0.10 0.10 0.00 0.05 0.15
7K 0.10 0.00 0.00 0.05 0.10
8K 0.15 0.15 0.10 0.20 0.20
Gemma3-12B-Instruct 1K 0.12 0.00 0.12 0.00 0.12
2K 0.25 0.25 0.00 0.12 0.12
3K 0.00 0.00 0.00 0.12 0.00
4K 0.00 0.00 0.00 0.11 0.00
5K 0.12 0.00 0.00 0.00 0.12
6K 0.00 0.00 0.00 0.00 0.14
7K 0.00 0.00 0.00 0.00 0.33
8K 0.00 0.00 0.00 0.00 0.17
Table 15: Comparison of FineSurE error rates for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest error rate in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.53 0.27 0.60 0.20 0.80
2K 0.80 0.53 0.67 0.80 1.47
3K 1.75 1.12 1.12 1.00 2.25
4K 0.89 1.22 1.44 0.44 2.44
5K 0.60 1.07 1.13 0.87 2.07
6K 0.83 0.67 0.67 0.67 1.50
7K 0.90 0.60 1.00 0.90 1.90
8K 0.67 0.92 0.75 0.83 1.17
Qwen2.5-7B-Instruct 1K 0.23 0.31 0.23 0.15 0.15
2K 0.31 0.00 0.54 0.38 0.23
3K 0.25 0.25 0.42 0.33 0.17
4K 0.50 0.20 0.20 0.50 0.70
5K 0.20 0.30 0.00 0.40 0.30
6K 0.25 0.92 0.67 0.67 0.33
7K 0.64 0.82 0.82 0.64 0.27
8K 0.10 0.10 0.30 0.30 1.2
GPT-4o mini 1K 0.20 0.25 0.10 0.20 0.15
2K 0.10 0.20 0.30 0.15 0.40
3K 0.43 0.36 0.57 0.21 0.64
4K 0.67 0.28 0.39 0.17 0.72
5K 0.31 0.50 0.25 0.12 0.31
6K 0.39 0.61 0.56 0.50 0.72
7K 0.50 0.71 0.36 0.36 0.50
8K 0.67 0.75 0.92 0.92 0.75
Gemma3-12B-Instruct 1K 0.00 0.00 0.00 0.00 0.20
2K 0.00 0.00 0.00 0.00 0.12
3K 0.30 0.30 0.00 0.30 0.30
4K 0.40 0.60 0.80 1.00 0.40
5K 0.25 0.00 0.25 0.25 0.75
6K 0.12 0.12 0.25 0.38 0.12
7K 0.00 0.00 0.00 0.00 0.14
8K 0.00 0.17 0.00 0.33 0.00
Table 16: Comparison of FineSurE error rates for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest error rate in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.66 0.63 0.63 0.62 0.57 6.5
2K 0.70 0.62 0.64 0.60 0.57 7.0
3K 0.70 0.64 0.63 0.63 0.61 4.0
4K 0.60 0.66 0.63 0.61 0.58 4.5
5K 0.70 0.65 0.66 0.64 0.61 5.25
6K 0.70 0.67 0.69 0.66 0.62 6.0
7K 0.74 0.67 0.64 0.65 0.60 7.5
8K 0.68 0.65 0.60 0.64 0.55 9.25
Qwen2.5-7B-Instruct 1K 0.70 0.58 0.65 0.61 0.58 5.5
2K 0.81 0.61 0.64 0.57 0.52 13.75
3K 0.70 0.62 0.57 0.60 0.54 8.25
4K 0.73 0.68 0.63 0.59 0.57 8.75
5K 0.71 0.65 0.66 0.57 0.57 7.75
6K 0.70 0.60 0.62 0.62 0.58 5.5
7K 0.72 0.57 0.61 0.61 0.54 8.75
8K 0.66 0.60 0.64 0.61 0.59 3.75
GPT-4o mini 1K 0.68 0.60 0.55 0.51 0.51 7.5
2K 0.73 0.58 0.58 0.57 0.50 11.5
3K 0.72 0.59 0.58 0.57 0.53 8.5
4K 0.68 0.60 0.55 0.56 0.52 7.75
5K 0.68 0.61 0.60 0.56 0.52 9.25
6K 0.67 0.61 0.60 0.57 0.52 9.25
7K 0.68 0.62 0.58 0.57 0.53 8.25
8K 0.69 0.62 0.56 0.56 0.49 11.75
Gemma3-12B-Instruct 1K 0.64 0.55 0.56 0.51 0.62 -5.5
2K 0.65 0.56 0.52 0.53 0.54 2.5
3K 0.65 0.63 0.55 0.55 0.53 6.5
4K 0.63 0.59 0.57 0.56 0.52 6.75
5K 0.61 0.57 0.62 0.57 0.53 6.25
6K 0.69 0.63 0.60 0.56 0.54 8.0
7K 0.68 0.61 0.63 0.67 0.56 8.75
8K 0.68 0.61 0.63 0.67 0.56 8.75
Table 17: Comparison of ROUGE-L based faithfulness scores and Sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.65 0.62 0.60 0.61 0.58 4.0
2K 0.65 0.66 0.64 0.62 0.57 7.25
3K 0.65 0.66 0.65 0.64 0.56 9.0
4K 0.62 0.61 0.60 0.64 0.60 1.75
5K 0.65 0.66 0.65 0.63 0.58 6.75
6K 0.65 0.71 0.71 0.69 0.61 8.0
7K 0.68 0.68 0.64 0.64 0.60 6.0
8K 0.62 0.67 0.67 0.67 0.62 3.75
Qwen2.5-7B-Instruct 1K 0.69 0.57 0.58 0.62 0.52 9.5
2K 0.62 0.62 0.60 0.60 0.54 7.0
3K 0.66 0.60 0.60 0.62 0.56 6.0
4K 0.63 0.61 0.61 0.66 0.57 5.75
5K 0.64 0.65 0.63 0.63 0.60 3.75
6K 0.65 0.63 0.64 0.61 0.57 6.25
7K 0.65 0.66 0.62 0.60 0.57 6.25
8K 0.64 0.66 0.63 0.61 0.57 6.5
GPT-4o mini 1K 0.69 0.57 0.55 0.58 0.52 7.75
2K 0.61 0.55 0.54 0.53 0.50 5.75
3K 0.59 0.55 0.55 0.50 0.47 7.75
4K 0.62 0.58 0.56 0.55 0.53 4.75
5K 0.59 0.55 0.55 0.52 0.46 9.25
6K 0.59 0.56 0.55 0.53 0.49 6.75
7K 0.58 0.56 0.53 0.51 0.49 5.5
8K 0.58 0.54 0.53 0.50 0.47 6.75
Gemma3-12B-Instruct 1K 0.66 0.57 0.60 0.61 0.47 14.0
2K 0.63 0.60 0.60 0.61 0.57 4.0
3K 0.59 0.59 0.54 0.57 0.56 1.25
4K 0.61 0.56 0.58 0.58 0.58 0.25
5K 0.62 0.60 0.61 0.63 0.54 7.5
6K 0.60 0.63 0.61 0.59 0.57 3.75
7K 0.65 0.65 0.60 0.61 0.57 5.75
8K 0.60 0.64 0.60 0.60 0.56 5.0
Table 18: Comparison of ROUGE-L based faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.81 0.79 0.85 0.82 0.79 2.75
2K 0.84 0.79 0.81 0.81 0.77 4.25
3K 0.87 0.85 0.84 0.83 0.79 5.75
4K 0.91 0.90 0.88 0.86 0.88 0.75
5K 0.85 0.85 0.86 0.83 0.83 1.75
6K 0.89 0.85 0.85 0.83 0.81 4.5
7K 0.91 0.81 0.85 0.87 0.81 5.0
8K 0.87 0.84 0.82 0.82 0.79 4.75
Arxiv 1K 0.71 0.79 0.72 0.80 0.69 6.5
2K 0.82 0.71 0.76 0.69 0.75 -0.5
3K 0.74 0.87 0.78 0.82 0.77 3.25
4K 0.87 0.86 0.84 0.85 0.79 6.5
5K 0.84 0.85 0.79 0.82 0.81 1.5
6K 0.82 0.83 0.77 0.85 0.90 -8.25
7K 0.86 0.82 0.89 0.84 0.79 6.25
8K 0.80 0.81 0.75 0.78 0.81 -2.5
Pubmed 1K 0.88 0.87 0.88 0.79 0.86 -0.5
2K 0.89 0.85 0.81 0.87 0.85 0.5
3K 0.91 0.89 0.86 0.91 0.90 -0.75
4K 0.93 0.90 0.92 0.87 0.92 -1.5
5K 0.91 0.84 0.89 0.91 0.85 3.75
6K 0.91 0.88 0.94 0.91 0.91 0.0
7K 0.91 0.89 0.92 0.91 0.92 -1.25
8K 0.91 0.92 0.89 0.87 0.90 -0.25
Govreport 1K 0.87 0.82 0.86 0.83 0.81 3.5
2K 0.94 0.88 0.90 0.93 0.93 -1.75
3K 0.91 0.92 0.90 0.91 0.91 0.0
4K 0.94 0.83 0.88 0.85 0.84 3.5
5K 0.94 0.93 0.94 0.93 0.93 0.5
6K 0.93 0.91 0.92 0.94 0.87 5.5
7K 0.91 0.89 0.94 0.97 0.89 3.75
8K 0.90 0.93 0.93 0.90 0.93 -1.5
Table 19: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.78 0.82 0.75 0.75 0.75 2.5
2K 0.82 0.79 0.75 0.76 0.73 5.0
3K 0.77 0.81 0.80 0.81 0.71 8.75
4K 0.79 0.80 0.73 0.77 0.74 3.25
5K 0.80 0.83 0.83 0.84 0.79 3.5
6K 0.80 0.82 0.77 0.83 0.72 8.5
7K 0.84 0.87 0.84 0.79 0.77 6.5
8K 0.84 0.85 0.82 0.82 0.77 6.25
Arxiv 1K 0.73 0.81 0.81 0.73 0.72 5.0
2K 0.80 0.75 0.78 0.72 0.75 1.25
3K 0.78 0.77 0.76 0.75 0.72 4.5
4K 0.74 0.80 0.70 0.78 0.74 1.5
5K 0.81 0.80 0.84 0.83 0.77 5.0
6K 0.70 0.77 0.72 0.72 0.74 -1.25
7K 0.86 0.81 0.76 0.74 0.71 8.25
8K 0.79 0.74 0.81 0.77 0.68 9.75
Pubmed 1K 0.89 0.87 0.93 0.87 0.89 0.0
2K 0.84 0.87 0.81 0.88 0.81 4.0
3K 0.86 0.89 0.90 0.90 0.87 1.75
4K 0.83 0.89 0.84 0.86 0.79 6.5
5K 0.87 0.87 0.86 0.85 0.85 1.25
6K 0.87 0.90 0.86 0.81 0.80 6.0
7K 0.91 0.89 0.87 0.82 0.82 5.25
8K 0.90 0.82 0.86 0.81 0.71 13.75
Govreport 1K 0.88 0.86 0.86 0.89 0.85 2.25
2K 0.89 0.90 0.88 0.91 0.91 -1.5
3K 0.87 0.92 0.91 0.91 0.88 2.25
4K 0.90 0.89 0.94 0.90 0.82 8.75
5K 0.94 0.91 0.90 0.91 0.89 2.5
6K 0.91 0.91 0.90 0.92 0.92 -1.0
7K 0.93 0.94 0.93 0.93 0.91 2.25
8K 0.92 0.93 0.95 0.92 0.85 8.0
Table 20: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.73 0.73 0.75 0.74 0.71 3.0
2K 0.77 0.76 0.75 0.73 0.73 2.25
3K 0.79 0.80 0.76 0.76 0.69 8.75
4K 0.81 0.82 0.81 0.79 0.76 4.75
5K 0.80 0.79 0.79 0.76 0.75 3.5
6K 0.81 0.80 0.80 0.76 0.71 8.25
7K 0.85 0.76 0.78 0.78 0.70 9.25
8K 0.80 0.76 0.73 0.74 0.63 12.75
Arxiv 1K 0.69 0.76 0.79 0.81 0.71 5.25
2K 0.75 0.76 0.80 0.76 0.70 6.75
3K 0.79 0.86 0.84 0.83 0.70 13.0
4K 0.79 0.83 0.83 0.83 0.75 7.0
5K 0.82 0.84 0.84 0.83 0.79 4.25
6K 0.81 0.83 0.81 0.81 0.81 0.5
7K 0.85 0.86 0.86 0.86 0.76 9.75
8K 0.78 0.83 0.83 0.81 0.77 4.25
Pubmed 1K 0.81 0.83 0.81 0.73 0.79 0.5
2K 0.82 0.84 0.78 0.81 0.74 7.25
3K 0.83 0.84 0.81 0.83 0.76 6.75
4K 0.81 0.82 0.78 0.80 0.76 4.25
5K 0.81 0.80 0.83 0.82 0.73 8.5
6K 0.79 0.80 0.86 0.83 0.78 4.0
7K 0.80 0.81 0.82 0.79 0.78 2.5
8K 0.83 0.84 0.81 0.84 0.75 8.0
Govreport 1K 0.72 0.79 0.80 0.82 0.76 2.25
2K 0.83 0.82 0.86 0.83 0.77 6.5
3K 0.83 0.84 0.83 0.82 0.81 2.0
4K 0.82 0.76 0.83 0.78 0.78 1.75
5K 0.83 0.81 0.86 0.82 0.82 1.0
6K 0.84 0.83 0.87 0.86 0.82 3.0
7K 0.83 0.86 0.89 0.89 0.79 7.75
8K 0.84 0.82 0.87 0.85 0.81 3.5
Table 21: Comparison of NLI model based faithfulness scores and Sensitivity for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.76 0.75 0.69 0.68 0.68 4.0
2K 0.73 0.76 0.73 0.70 0.63 10.0
3K 0.71 0.72 0.72 0.74 0.62 10.25
4K 0.71 0.75 0.70 0.70 0.67 4.5
5K 0.71 0.75 0.79 0.79 0.70 6.0
6K 0.69 0.76 0.73 0.76 0.61 12.5
7K 0.78 0.79 0.75 0.76 0.67 10.0
8K 0.74 0.79 0.74 0.77 0.67 9.0
Arxiv 1K 0.77 0.82 0.81 0.80 0.74 6.0
2K 0.76 0.78 0.80 0.72 0.71 5.5
3K 0.75 0.79 0.81 0.78 0.68 10.25
4K 0.71 0.81 0.74 0.77 0.73 2.75
5K 0.77 0.83 0.85 0.81 0.74 7.5
6K 0.69 0.82 0.79 0.75 0.75 1.25
7K 0.86 0.86 0.80 0.80 0.77 6.0
8K 0.79 0.81 0.85 0.82 0.73 8.75
Pubmed 1K 0.75 0.85 0.84 0.80 0.78 3.0
2K 0.76 0.83 0.82 0.83 0.72 9.0
3K 0.81 0.85 0.86 0.78 0.76 6.5
4K 0.77 0.82 0.79 0.79 0.68 11.25
5K 0.78 0.81 0.81 0.78 0.70 9.5
6K 0.80 0.87 0.83 0.77 0.74 7.75
7K 0.85 0.85 0.84 0.79 0.77 6.25
8K 0.80 0.79 0.81 0.77 0.67 12.25
Govreport 1K 0.80 0.80 0.85 0.87 0.78 5.0
2K 0.82 0.85 0.83 0.85 0.78 5.75
3K 0.81 0.86 0.89 0.81 0.82 2.25
4K 0.76 0.86 0.86 0.82 0.77 5.5
5K 0.82 0.85 0.88 0.82 0.78 6.25
6K 0.86 0.88 0.88 0.87 0.80 7.25
7K 0.86 0.91 0.89 0.88 0.83 5.5
8K 0.84 0.87 0.89 0.84 0.80 6.0
Table 22: Comparison of NLI model based faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.